# Categorical Flow Matching on Statistical Manifolds

### v2

@@ A.1 Riemannian Manifold
A Riemannian manifold $\mathcal{M}$ is a real, smooth manifold equipped with a positive definite inner product $g$ on the tangent space $T_{x}(\mathcal{M})$ at each point $x\in\mathcal{M}$. Let $T\mathcal{M}=\bigcup_{x\in\mathcal{M}}T_{x}(\mathcal{M})$ be the tangent bundle of the manifold $\mathcal{M}$, a time-dependent vector field on $\mathcal{M}$ is a mapping $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ where $u_{t}(x)\in T_{x}(\mathcal{M})$. A geodesic is a locally distance-minimizing curve on the manifold. The existence and the uniqueness of the geodesic state that for any point $x\in\mathcal{M}$ and for any tangent vector $u\in T_{x}(\mathcal{M})$, there exists a unique geodesic $\gamma:[0,1]\to\mathcal{M}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=u$. The exponential map $\exp:\mathcal{M}\times T\mathcal{M}\to\mathcal{M}$ is uniquely defined to be $\exp_{x}(u):=\gamma(1)$. The logarithm map $\log:\mathcal{M}\times\mathcal{M}\to T\mathcal{M}$ is defined as the inverse mapping of the exponential map such that $\exp_{x}(\log_{x}(y))\equiv y,\forall x,y\in\mathcal{M}$***Note that the formulation here is slightly different from Sec.2.1 where we fix a point $\mu$ on $\mathcal{P}$.. With the exponential map and logarithm map, the time-dependent flow can be compactly written as time interpolation along the geodesic:
$x_{t}:=\psi_{t}(x_{t}|x_{0},x_{1})=\exp_{x_{0}}(t\log_{x_{0}}x_{1}),\quad t\in[0,1]$ (15)
@@
$\frac{d^{2}x^{k}}{dt^{2}}+\Gamma_{ij}^{k}\frac{dx^{i}}{dt}\frac{dx^{j}}{dt}=0,\quad k=1,2,\dots,n$ (16)
@@
$\Gamma_{ij}^{k}=\frac{1}{2}g^{km}\left(\frac{\partial g_{mi}}{\partial x^{j}}+\frac{\partial g_{mj}}{\partial x^{i}}-\frac{\partial g_{ij}}{\partial x^{m}}\right),\quad i,j,k=1,2,\dots,n$ (17)
@@ A.2 Spherical Manifold
The positive orthant of the unit $(n-1)$-sphere $S^{n-1}_{+}$ is a $(n-1)$-dimensional manifold. The sphere can be embedded into the ambient Euclidean manifold $\mathbb{R}^{n}$ so it inherits the canonical inner product as
$\langle u,v\rangle_{S}=\langle u,v\rangle=\sum_{i=1}^{n}u_{i}v_{i},\quad u,v\in T_{x}(S^{n-1}_{+})$ (18)
@@
$\log_{x}(y)=\arccos(\langle x,y\rangle)\frac{P_{x}(y-x)}{\|P_{x}(y-x)\|_{2}}$ (20)
@@ A.3 Statistical Manifold of Categorical Distributions
$g_{jk}(\mu)=\frac{\delta_{jk}}{\mu_{j}}+\frac{1}{\mu_{n}},\quad 1\leq i,j,\leq n-1$ (21)
@@
$\left\langle\frac{\partial\pi(\mu)}{\partial u},\frac{\partial\pi(\mu)}{\partial v}\right\rangle=\frac{1}{4}\langle u,v\rangle_{\mu},\quad u,v\in T_{\mu}(\mathcal{P})$ (22)
@@
$\left\langle\frac{\partial\pi(\mu)}{\partial u},\frac{\partial\pi(\mu)}{\partial v}\right\rangle=\left\langle\left.\frac{d}{dt}\pi(\mu+tu)\right|_{t=0},\left.\frac{d}{dt}\pi(\mu+tv)\right|_{t=0}\right\rangle=\sum_{i=1}^{n}\frac{u_{i}}{2\sqrt{\mu_{i}}}\frac{v_{i}}{2\sqrt{\mu_{i}}}=\frac{1}{4}\langle u,v\rangle_{\mu}$
Note that the infinitesimal arc length on the geodesic can be expressed as $ds^{2}=\|dx\|_{g}^{2}=\sum_{jk}dx_{j}g_{jk}(x)dx_{k}$ where $\|\cdot\|_{g}$ is the canonical Riemannian norm induced by the inner product. Integrating over the geodesic, we can easily arrive at the result in Proposition 1. Based on this, the exponential map and logarithm map on this statistical manifold can be written as
$\exp_{\mu}(u)$ $=\left(\sqrt{\mu}\cos\left\|\frac{u}{2\sqrt{\mu}}\right\|_{2}+\frac{u}{2\sqrt{\mu}}\operatorname{sinc}\left\|\frac{u}{2\sqrt{\mu}}\right\|_{2}\right)^{2}$ (23)
$\log_{\mu}(\nu)$ $=\frac{2\arccos(\langle\sqrt{\mu},\sqrt{\nu}\rangle)}{\sqrt{1-\langle\sqrt{\mu},\sqrt{\nu}\rangle}}(\sqrt{\mu\odot\nu}-\langle\sqrt{\mu},\sqrt{\nu}\rangle\mu)$ (24)
@@ A.4 Other Statistical Manifolds
$d_{\text{mul}}(\mu,\nu)=\sqrt{n}d_{\text{cat}}(\mu,\nu)=2\sqrt{n}\arccos\left(\sum_{i=1}^{n}\sqrt{\mu_{i}\nu_{i}}\right)$ (27)
@@ B.1 Negative Log-Likelihood (NLL) Calculation
$p(\theta)=\frac{1}{\pi\sqrt{\theta(1-\theta)}},\quad\theta\in(0,1)$ (28)
@@
$=\mathbb{E}_{q(\mu|\delta)}\left[\log p(\delta|\mu)+\log p(\mu)-\log q(\mu|\delta)\right]$
@@
$\tilde{\theta}_{i}\sim\mathrm{Exp}(1),\theta_{i}=\tilde{\theta}_{i}/\sum_{k=1}^{n}\tilde{\theta}_{k},\quad i=1,2,\dots,n$ (33)
@@
$\tilde{x}_{n-1}$ $=\sin^{2}\varphi_{1}\dots\sin^{2}\varphi_{n-2}\cos^{2}\varphi_{n-1}$
@@
$\frac{d}{dt}\left[\begin{array}[]{c}x_{t}\\ \log p^{\text{ODE}}_{t}\end{array}\right]=\left[\begin{array}[]{c}v_{t}(x_{t})\\ -\operatorname{div}_{g}(v_{t})(x_{t})\end{array}\right]$ (41)
@@
$\mathrm{div}\>v_{t}(x)=\mathbb{E}_{\varepsilon\sim\mathcal{N}(0,I)}[\varepsilon^{\top}\nabla v_{t}(x)\varepsilon]$ (42)
@@
2:Apply the diffeomorphism in Eq.(5) to obtain $\tilde{x}_{1}=\pi(\tilde{\mu}_{1})$ and calculate $\log|\det d\pi^{-1}(\tilde{x}_{1})|$.
@@
4:Apply $\pi^{-1}$ to obtain $\mu_{0}=\pi^{-1}(x_{0})$ and calculate $\log|\det d\pi(\mu_{0})|$.
@@ B.2 Bits-Per-Character (BPC) Calculation
[…]
@@
[…]
@@ B.3 GPT-J-6B Likelihood
$\mathcal{L}_{\text{GPT}}=-\frac{1}{K}\sum_{k=1}^{K}\log p\left(w_{k}|w_{1:k-1}\right)$ (46)
@@ C.1 Model Parameterization
$v_{t}(x_{t})=\tilde{v}_{t}(x_{t})-\langle x_{t},\tilde{v}_{t}(x_{t})\rangle x_{t}$ (47)
@@
$v_{t}(x_{t})=\tilde{v}_{t}(x_{t})-\frac{1}{n}\sum_{i=1}^{n}\tilde{v}_{t}(x_{t})_{i}$ (48)
@@
8:     Calculate the conditional vector field $u^{S}_{t}(x_{t}|x_{0},x_{1})=\frac{d}{dt}x_{t}=\log_{x_{t}}(x_{1})/(1-t)$.

### v3

@@ A.1 Riemannian Manifold
A Riemannian manifold $\mathcal{M}$ is a real, smooth manifold equipped with a positive definite inner product $g$ on the tangent space $T_{x}(\mathcal{M})$ at each point $x\in\mathcal{M}$. Let $T\mathcal{M}=\bigcup_{x\in\mathcal{M}}T_{x}(\mathcal{M})$ be the tangent bundle of the manifold $\mathcal{M}$, a time-dependent vector field on $\mathcal{M}$ is a mapping $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ where $u_{t}(x)\in T_{x}(\mathcal{M})$. A geodesic is a locally distance-minimizing curve on the manifold. The existence and the uniqueness of the geodesic state that for any point $x\in\mathcal{M}$ and for any tangent vector $u\in T_{x}(\mathcal{M})$, there exists a unique geodesic $\gamma:[0,1]\to\mathcal{M}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=u$. The exponential map $\exp:\mathcal{M}\times T\mathcal{M}\to\mathcal{M}$ is uniquely defined to be $\exp_{x}(u):=\gamma(1)$. The logarithm map $\log:\mathcal{M}\times\mathcal{M}\to T\mathcal{M}$ is defined as the inverse mapping of the exponential map such that $\exp_{x}(\log_{x}(y))\equiv y,\forall x,y\in\mathcal{M}$. Note that the formulation here is slightly different from Sec.2.1 where we fix a point $\mu$ on $\mathcal{P}$. With the exponential map and logarithm map, the time-dependent flow can be compactly written as time interpolation along the geodesic:
$x_{t}:=\psi_{t}(x_{t}|x_{0},x_{1})=\exp_{x_{0}}(t\log_{x_{0}}x_{1}),\quad t\in[0,1].$ (15)
@@
$\frac{\mathop{}\!\mathrm{d}^{2}x^{k}}{\mathop{}\!\mathrm{d}t^{2}}+\Gamma_{ij}^{k}\frac{\mathop{}\!\mathrm{d}x^{i}}{\mathop{}\!\mathrm{d}t}\frac{\mathop{}\!\mathrm{d}x^{j}}{\mathop{}\!\mathrm{d}t}=0,\quad k=1,2,\dots,n,$ (16)
@@
$\Gamma_{ij}^{k}=\frac{1}{2}g^{km}\left(\frac{\partial g_{mi}}{\partial x^{j}}+\frac{\partial g_{mj}}{\partial x^{i}}-\frac{\partial g_{ij}}{\partial x^{m}}\right),\quad i,j,k=1,2,\dots,n,$ (17)
@@ A.2 Spherical Manifold
The positive orthant of the unit $(n-1)$-sphere $S^{n-1}_{+}$ is a $(n-1)$-dimensional manifold. The sphere can be embedded into the ambient Euclidean manifold $\mathbb{R}^{n}$ such that it inherits the canonical inner product as
$\langle u,v\rangle_{S}=\langle u,v\rangle=\sum_{i=1}^{n}u_{i}v_{i},\quad u,v\in T_{x}(S^{n-1}_{+}).$ (18)
@@
$\log_{x}(y)=\arccos(\langle x,y\rangle)\frac{P_{x}(y-x)}{\|P_{x}(y-x)\|_{2}},$ (20)
@@ A.3 Statistical Manifold of Categorical Distributions
$g_{jk}(\mu)=\frac{\delta_{jk}}{\mu_{j}}+\frac{1}{\mu_{n}},\quad 1\leq i,j,\leq n-1,$ (21)
@@
$\left\langle\frac{\partial\pi(\mu)}{\partial u},\frac{\partial\pi(\mu)}{\partial v}\right\rangle=\frac{1}{4}\langle u,v\rangle_{\mu},\quad u,v\in T_{\mu}(\mathcal{P}).$ (22)
@@
$\left\langle\frac{\partial\pi(\mu)}{\partial u},\frac{\partial\pi(\mu)}{\partial v}\right\rangle=\left\langle\left.\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\pi(\mu+tu)\right|_{t=0},\left.\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\pi(\mu+tv)\right|_{t=0}\right\rangle=\sum_{i=1}^{n}\frac{u_{i}}{2\sqrt{\mu_{i}}}\frac{v_{i}}{2\sqrt{\mu_{i}}}=\frac{1}{4}\langle u,v\rangle_{\mu}.$
Note that the infinitesimal arc length on the geodesic can be expressed as $\mathop{}\!\mathrm{d}s^{2}=\|\mathop{}\!\mathrm{d}x\|_{g}^{2}=\sum_{jk}\mathop{}\!\mathrm{d}x_{j}g_{jk}(x)\mathop{}\!\mathrm{d}x_{k}$ where $\|\cdot\|_{g}$ is the canonical Riemannian norm induced by the inner product. Integrating over the geodesic, we can easily arrive at the result in Proposition 1. Based on this, the exponential map and logarithm map on this statistical manifold can be written as
$\exp_{\mu}(u)$ $=\left(\sqrt{\mu}\cos\left\|\frac{u}{2\sqrt{\mu}}\right\|_{2}+\frac{u}{2\sqrt{\mu}}\operatorname{sinc}\left\|\frac{u}{2\sqrt{\mu}}\right\|_{2}\right)^{2},$ (23)
$\log_{\mu}(\nu)$ $=\frac{2\arccos(\langle\sqrt{\mu},\sqrt{\nu}\rangle)}{\sqrt{1-\langle\sqrt{\mu},\sqrt{\nu}\rangle}}(\sqrt{\mu\odot\nu}-\langle\sqrt{\mu},\sqrt{\nu}\rangle\mu),$ (24)
@@ A.4 Other Statistical Manifolds
$d_{\text{mul}}(\mu,\nu)=\sqrt{n}d_{\text{cat}}(\mu,\nu)=2\sqrt{n}\arccos\left(\sum_{i=1}^{n}\sqrt{\mu_{i}\nu_{i}}\right).$ (27)
@@ B.1 Negative Log-Likelihood (NLL) Calculation
$p(\theta)=\frac{1}{\pi\sqrt{\theta(1-\theta)}},\quad\theta\in(0,1).$ (28)
@@
$=\mathbb{E}_{q(\mu|\delta)}\left[\log p(\delta|\mu)+\log p(\mu)-\log q(\mu|\delta)\right].$
@@
$\tilde{\theta}_{i}\sim\mathrm{Exp}(1),\theta_{i}=\tilde{\theta}_{i}/\sum_{k=1}^{n}\tilde{\theta}_{k},\quad i=1,2,\dots,n.$ (33)
@@
$\tilde{x}_{n-1}$ $=\sin^{2}\varphi_{1}\dots\sin^{2}\varphi_{n-2}\cos^{2}\varphi_{n-1}.$
@@
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\left[\begin{array}[]{c}x_{t}\\ \log p^{\text{ODE}}_{t}\end{array}\right]=\left[\begin{array}[]{c}v_{t}(x_{t})\\ -\operatorname{div}_{g}(v_{t})(x_{t})\end{array}\right].$ (41)
@@
$\mathrm{div}\>v_{t}(x)=\mathbb{E}_{\varepsilon\sim\mathcal{N}(0,I)}[\varepsilon^{\top}\nabla v_{t}(x)\varepsilon].$ (42)
@@
2:Apply the diffeomorphism in Eq.(5) to obtain $\tilde{x}_{1}=\pi(\tilde{\mu}_{1})$ and calculate $\log|\det\mathop{}\!\mathrm{d}\pi^{-1}(\tilde{x}_{1})|$.
@@
4:Apply $\pi^{-1}$ to obtain $\mu_{0}=\pi^{-1}(x_{0})$ and calculate $\log|\det\mathop{}\!\mathrm{d}\pi(\mu_{0})|$.
@@ B.2 Bits-Per-Character (BPC) Calculation
$\mathcal{L}_{\text{ELBO}}=-\mathbb{E}_{\mu_{1}\sim q(\mu)}\left[\int_{0}^{1}\frac{1}{1-t}\log\langle\hat{\mu}_{1}(\mu_{t}),\mu_{1}\rangle\mathop{}\!\mathrm{d}t\right]$ (43)
@@
$\mathcal{L}_{\text{ELBO}}=-\mathbb{E}_{\mu_{1}\sim q(\mu)}\left[\int_{0}^{\infty}\log\langle\hat{\mu}_{1}(\mu_{t}),\mu_{1}\rangle\mathop{}\!\mathrm{d}s\right]$ (45)
@@ B.3 GPT-J-6B Likelihood
$\mathcal{L}_{\text{GPT}}=-\frac{1}{K}\sum_{k=1}^{K}\log p\left(w_{k}|w_{1:k-1}\right),$ (46)
@@ C.1 Model Parameterization
$v_{t}(x_{t})=\tilde{v}_{t}(x_{t})-\langle x_{t},\tilde{v}_{t}(x_{t})\rangle x_{t}.$ (47)
@@
$v_{t}(x_{t})=\tilde{v}_{t}(x_{t})-\frac{1}{n}\sum_{i=1}^{n}\tilde{v}_{t}(x_{t})_{i}.$ (48)
@@
8:     Calculate the conditional vector field $u^{S}_{t}(x_{t}|x_{0},x_{1})=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}x_{t}=\log_{x_{t}}(x_{1})/(1-t)$.